OpenStudy (anonymous):

Which graph represents the function of f(x) = the quantity of 9 x squared plus 9 x minus 18, all over 3 x plus 6?

OpenStudy (whimsical):

\[\frac{ 9x^2+9x-18 }{ 3x+6 }\] is it this equation?

OpenStudy (anonymous):

yes @Whimsical

OpenStudy (whimsical):

\[\frac{ 9(x+2)(x-1) }{ 3(x+2) }\] simplified the equation what do you get?

OpenStudy (anonymous):

3(x-1) ?

OpenStudy (whimsical):

yes and what kind of graph is that?

OpenStudy (whimsical):

are you doing okay or do you need more hints?

OpenStudy (anonymous):

is it a linear equation?

OpenStudy (anonymous):

i'm not sure what kind of graph it is

OpenStudy (whimsical):

yes, can you tell me what makes a straight line graph

OpenStudy (anonymous):

the coordinates can be easily placed with rise over run after the y-intercept is placed

OpenStudy (whimsical):

yes so the equations is: y=mx+c m= gradient c= y-intercept can you find m and c? i got to go now, hope u can find the straight line graph equation,

OpenStudy (whimsical):

btw 3(x-1)=mx+c in case you need more help

OpenStudy (anonymous):

m= 3x y = -3

OpenStudy (whimsical):

u can sketch the graph now  and m = 3 not 3x

OpenStudy (anonymous):

oh, okay! I get it now! Thank you so much!

OpenStudy (trojanpoem):

f(-2) = 4*9 - 18 - 18/ 6- 6 = 0/0 (undefined) Your function is undefined as x = -2, so you simply draw an asymptote at that point. and the intercept of the asymptote and your line is a gap. See this: http://www.purplemath.com/modules/grphrtnl4.htm

OpenStudy (welshfella):

sometimes this is called a 'hole' in the graph